OpenStudy (anonymous):

Calc 2, representing functions as power series question.

OpenStudy (anonymous):

Evaluate the indefinite integral as a power series and find the radius of convergence.\[\int\limits_{}^{}\frac{x-\tan^{-1} x}{x^3}\]

OpenStudy (anonymous):

I know what I want to do is get it into the form \[\frac{1}{1-r}\] so I can relate it to a geometric series but honestly not even sure where to start.

OpenStudy (anonymous):

are you sure about the \(\frac{1}{1-r}\) part?

OpenStudy (anonymous):

\[\arctan(x)=x-\frac{x^3}{3}+\frac{x^5}{5}-\frac{x^7}{7}+...\]

OpenStudy (anonymous):

How else could I relate it to a geometric series? (If there is another way I could express it, then all we've been taught so far is to relate it to a geometric series, which would be the 1/1-r)

OpenStudy (anonymous):

\[x-\arctan(x)=\frac{x^3}{3}-\frac{x^5}{5}+\frac{x^7}{7}...\]

OpenStudy (anonymous):

divide by \(x^2\) to get your power series

OpenStudy (anonymous):

relating to a geometric series is a nice way to do it as a gimmick, if you have something annoying to compute

OpenStudy (anonymous):

but in this case, assuming you know the power series for arctangent, then you can compute it directly

OpenStudy (anonymous):

Do not know the power series for arctangent. and I'm getting it to be \[\sum_{n=0}^{\infty}\frac{-x^{2n-1}}{n}\]And that definitely isn't what you had it expressed as.. :/

OpenStudy (anonymous):

oh wait. I think I know what I did wrong. um, well not quite. \[\sum_{n=0}^{\infty}\frac{-x^{2n+1}}{n+1}\]

OpenStudy (anonymous):

yes it is except you need \[\sum_{n=0}^{\infty}\frac{(-1)^nx^{2n+1}}{n}\]

OpenStudy (anonymous):

that you can find the by "geometric series" trick

OpenStudy (anonymous):

the derivative of \(\arctan(x)\) is \(\frac{1}{1+x^2}\) which you can rewrite as \[\frac{1}{1-(-x^2)}=\sum_{n=0}^{\infty}(-1)^nx^{2n}\] and then

OpenStudy (anonymous):

then integrate term by term

OpenStudy (anonymous):

so i guess that is the part of this problem that uses geometric series the rest is more or less algebra

OpenStudy (anonymous):

okay. but then, shouldn't the power x is raised to be greater than the denominator? Since the power would be 2n+1 and the denominator would just be n+1?

OpenStudy (anonymous):

i had a mistake above, sorry it should be \[\sum_{n=0}^{\infty}\frac{(-1)^nx^{2n+1}}{2n+1}\]

OpenStudy (anonymous):

OH! Okay, I see it now.

OpenStudy (anonymous):

it is actually sort of easier to do the algebra without sigma notation

OpenStudy (anonymous):

Okay, now the algebra part..you can take out the n=0 term since that is just x and x is considered a constant right? and that gets you x-arctan(x). So the power series should be \[\Large\frac{x-\sum_{n=0}^{\infty}-1^n\frac{x^{2n+1}}{2n+1}}{x^3}\]? Or should that be n=1 and not n=0?

OpenStudy (anonymous):

i would do it without sigma notation, so see what the algebra gives you

OpenStudy (anonymous):

\[\arctan(x)=x-\frac{x^3}{3}+\frac{x^5}{5}-\frac{x^7}{7}+...\] \[x-\arctan(x)=\frac{x^3}{3}-\frac{x^5}{5}+\frac{x^7}{7}...\] \[\frac{x-\arctan(x)}{x^2}=\frac{x}{3}-\frac{x^2}{5}+\frac{x^5}{7}-...\]

OpenStudy (anonymous):

damn, another mistake \[\frac{x-\arctan(x)}{x^2}=\frac{x}{3}-\frac{x^3}{5}+\frac{x^5}{7}-...\]

OpenStudy (anonymous):

if it is easier for you to work with the notation, by all means do it, but i am kind of boneheaded that way. i like to see what exactly i get at each step

OpenStudy (anonymous):

so that would be \[x-arctan(x) = x-x-\frac{x^{3}}{3}+\frac{x^{5}}{5}-\frac{x^{7}}{7}...\] so divide by x-cubed...\[\frac{x-arctan(x)}{x^3} = -\frac{1}{3}+\frac{x^{2}}{5}-\frac{x^{4}}{7}...\]

OpenStudy (anonymous):

And I'll need it in sigma notation to answer it. So there's that.. lol

OpenStudy (anonymous):

oh you are dividing by \(x^3\) i am full or errors today

OpenStudy (anonymous):

but don't forget the distributive law!!!

OpenStudy (anonymous):

It's okay. I understood. And that's fewer errors than me ha XD and...oh, okay. Now I see.

OpenStudy (anonymous):

\[x-\arctan(x) = x-x-\frac{x^{3}}{3}+\frac{x^{5}}{5}-\frac{x^{7}}{7}...\] is wrong!

OpenStudy (anonymous):

\[x-\arctan(x) = x-(x-\frac{x^{3}}{3}+\frac{x^{5}}{5}-\frac{x^{7}}{7}...\]

OpenStudy (anonymous):

\[\frac{x-arctan(x)}{x^3} = \frac{1}{3}-\frac{x^{2}}{5}+\frac{x^{4}}{7}...\]

OpenStudy (anonymous):

that looks better now integrate term by term THEN rewrite in sigma notation at least that is how i would get it

OpenStudy (anonymous):

And so, we found arctan(x). and okay I got it. Thanks!

OpenStudy (anonymous):

yw